Q2d 400 Gen(s+4)(s2+4s+5)(s+4s+2-3s+2+3) Find the partial fraction expansion of F(s) and then use the Laplace transform tables to find f(t) ft)- cos( t+ oju(t) COS
Answer #1

Ans)

400 F(s) =

A=\lim_{s\rightarrow -4}(s+4)F(s)=\lim_{s\rightarrow -4}(s+4)\frac{400}{(s+4)(s^{2}+4s+5)}=80

A=80

F(s)=\frac{400}{(s+4)(s^{2}+4s+5)}=\frac{A(s^{2}+4s+5)+(Bs+C)(s+4)}{(s+4)(s^{2}+4s+5)}

Equating numerators we get

400=A(s^{2}+4s+5)+(Bs+C)(s+4)

400=As^{2}+4As+5A+Bs^{2}+Cs+4Bs+4C

400=(A+B)s^{2}+(4A+4B+C)s+5A+4C

Equating the coefficients

A+B=0,\: \: 4A+4B+C=0,\: \: 5A+4C=400

4(A+B)+C=0

4.0+C=0

C=0

5A+4C=400

5A+4*0=400

A=\frac{400}{5}=80

A=80

A+B=0\rightarrow B=-A=-80

B=-80

F(s)=\frac{400}{(s+4)(s^{2}+4s+5)}=\frac{80}{s+4}+\frac{-80s+0}{s^{2}+4s+5}

F(s)=\frac{80}{s+4}-\frac{80s}{s^{2}+4s+5}

F(s)=\frac{80}{s+4}-80(\frac{s+2-2}{(s+2)^{2}+1})

F(s)=\frac{80}{s+4}-80(\frac{s+2}{(s+2)^{2}+1}-\frac{2}{(s+2)^{2}+1})

Now taking inverse laplace transform we get

f(t)=80e^{-4t}-80(e^{-2t}cos(t)-2e^{-2t}sin(t))

\mathbf{f(t)=80e^{-4t}-80e^{-2t}(cos(t)-2sin(t))}